You are given the temperature T of an object in one of Celsius, Fahrenheit, and Kelvin scales.
Write a program to print T in all scales viz Celsius, Fahrenheit, and Kelvin.
Formula to convert from Fahrenheit F to Celsius C is C = (F - 32) * 5 / 9.
Formula to convert from Kelvin K to Celsius C is C = K - 273.
Here "C", "F", "K" represent that the temperature scale is in Celsius, Fahrenheit and Kelvin scales respectively.
The input contains the temperature (a number) and the unit of the temperature scale (C, F, K) without any space.
The output contains temperature in Celsius, Fahrenheit and Kelvin scales in each line in the format similar to input and the value of the temperature is rounded to 2 decimal places.
Input

The first line of the input contain a temperature Value in one of Celsius, Fahrenheit, and Kelvin scales.
Output

The first line of output should contain the Celsius value and the unit of the Celsius without any space.
The second line of output should contain the Fahrenheit value and the unit of the Fahrenheit without any space.
The third line of output should contain the Kelvin value and the unit of the Kelvin without any space.
Explanation

For example, if the given temperature Value is 25C then Celsius value is 25.0C, Fahrenheit value is 77.0F, and Kelvin value is 298.0K.
Sample Input 1
25C
Sample Output 1
25.0C
77.0F
298.0K

Sample Input 2
37.5F
Sample Output 2
3.06C
37.5F
276.06K
1
Expert's answer
2021-03-07T18:43:47-0500
T = input()
try:
    scale = T[-1:]
    temp = float(T[:-1])
except:
    print('incorrect input')
if scale == 'C':
    C = temp
    F = (C*9/5) + 32
    K = C + 273
elif scale == 'K':
    K = temp
    C = K - 273
    F = (C*9/5) + 32
elif scale == 'F':
    F = temp
    C = (F - 32)*5/9
    K = C + 273
else:
    print('incorrect input')


out_str = [str(round(C, 2)) + 'C',
           str(round(F, 2)) + 'F',
           str(round(K, 2)) + 'K']
print(*out_str, sep='\n')